OpenStudy (anonymous):

cube root of ((a^6)(b^9))/-64

OpenStudy (kinggeorge):

\[\Large \sqrt[3]{a^6b^9 \over -64}={\sqrt[3]{a^6b^9} \over \sqrt[3]{-64}}\]Can you tell me what the cube root of \(a^6b^9\) is?

OpenStudy (anonymous):

dont you do \[(a ^{6}b ^{9})^{1/3}\]

OpenStudy (kinggeorge):

It's the same exact thing, just a slightly different way of writing it. Can you solve that yourself?

OpenStudy (anonymous):

no i'm not sure how to. what do you do with the exponents in the numerator?

OpenStudy (kinggeorge):

In this case, you can do the following. \[\Large (a^6b^9)^{1/3}=a^{6/3}\cdot b^{9/3}=a^2b^3\]You can actually distribute the exponent to the terms inside the parentheses.

OpenStudy (kinggeorge):

So that would be the numerator. Could you tell me what \((-64)^{1/3}\) is?

OpenStudy (anonymous):

-4. thanks

OpenStudy (kinggeorge):

Exactly. You're welcome.